Assume the following ratios are constant:

Total asset turnover 3
Profit margin 5.9 %
Equity multiplier 1.5
Payout ratio 35 %


What is the sustainable growth rate? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
  
Sustainable growth rate            %
Answer #1

ROE=Profit margin*Total asset turnover*Equity multiplier

=(0.059*1.5*3)=0.2655

Retention ratio=1-payout ratio

=(1-0.35)=0.65

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.2655*0.65)/[1-(0.2655*0.65)]

=20.86%(Approx).